Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
8.5
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
8.5

File 1156, a ThAr lamp, looks like a BIAS. System got 2 instead of one calibration, first OK, second was not. Underline: note that 1156 is a bad CALIBRATION.
Verified before start of night ops: south west pump still like-leak noise; its attached manometer reports 100 PSI. The other 2 pumps, no like-leak noises, report 0 PSI.
Relative humidity in range 45% to 80%. Looks clear but likely hazy and with thin clouds. T-dewPoint reached a 3.2C difference.
DIMM2 monitor down at 05:48 UT. Pachon DIMM already closed. No data for last quarter of night.
While closing noted the DOME shutter mechanism sounds a bit strange. Could you please check it as well as the south west pump? Thanks.